Find the sum of all counting numbers less than 100 which are not multiples of 4.
Answer by edjones(8007)

Sn = n/2 (a[1] + a[n]) Sum of a finite series
.
n=99 for numbers up less than 100.
a[1}=1
.
99/2 (1+99)
=99*50
=4950
.
sum of multiples of 4 including 96.
n=24
a[1]=4
.
24/2 (4+96)
=24*50
=1200
.
4950-1200
=3,750
.
Ed